# Channels of Stellar-mass Black Hole Formation

### v1

@@ 4 Black Hole Formation Channels
The different primary stellar-mass black hole formation pathways we discuss in this paper are represented by four distinct channels. For each channel we provide example models: Channel 1: asymmetric, high-compactness vigorously exploding solar-metallicity 40-$M_{\odot}$ and 19.56-$M_{\odot}$ models (§4.1) that within seconds of bounce form lower-gap black holes, but later due to subsequent fallback leave a spectrum of black hole masses from the lower gap upwards; Channel 2: a weakly exploding, mildly asymmetric, solar-metallicity 23-$M_{\odot}$ model (§4.2) that experiences late-time fallback accretion and black hole formation on $\sim$1000-second timescales.; Channel 3: a 100-$M_{\odot}$ tenth-solar model (§4.3) that experiences a PPISN and then an aborted highly-asymmetrical supernova explosion that initially sends out a shock wave to thousands of kilometers which weakens significantly into a sound wave. Within hundreds of milliseconds, its PNS collapses into a black hole and within tens of seconds that black hole fattens to $\sim$37 $M_{\odot}$; and Channel 4: two models (12.25- and 14-$M_{\odot}$) (§4.4) that never reignited their stalled shock wave and would accrete over minutes to hours timescales until “quiescent” black hole formation. This is the oft-assumed “silent” channel. That this progenitor mass interval might evolve through this channel is quite unexpected and may be wrong. However, using the Sukhbold et al. (2018) progenitor suite, this is what we and others (Couch et al., 2020; Pan et al., 2021; Sykes & Müller, 2024b) currently conclude. We do not discuss the secondary black hole formation channels through accretion onto a neutron star from a companion star, the merger of two neutron stars, or the late-time developments of a Thorne-Zytkov object (Thorne & Zytkow, 1977), all potentially important pathways, but outside the CCSN context.
Figure 4: Snapshots of an x-y slice in entropy space (per baryon per Boltzmann’s constant) of the inner region ($\pm$1000 km) of the explosion of the solar-metallicity 19.56-$M_{\odot}$ progenitor. These stills are taken at 1.0, 2.0, 3.0, and 3.767 seconds after bounce, the latter just before black hole formation. They reveal an asymmetric quasi-jetlike structure flapping about in response to the variations in the infalling plumes that are dancing over the PNS. The red regions are high-entropy neutrino-driven ejected, while the blue regions are mostly these lower-entropy (and more dense) infalling streams. Note the blue and red regions are “complementary” -- the driven ejecta avoid the infalling fingers. The degree of net asymmetry is correlated with the magnitude of the resultant recoil kick, with the corresponding neutrino emission anisotropy a small contributing factor. The 19.56-$M_{\odot}$ model explodes vigorously and asymmetrically and leaves behind a lower “gap” black hole. See text for discussion.
@@
Interestingly, if this 40-$M_{\odot}$ model had experienced binary mass transfer and been stripped to a “stripped-envelope” structure, the fallback mass would have been minimal, the ejected 56Ni mass would have been more significant, the total ejected mass would have been much lower, as would the residual black hole mass, and the kick to the black hole would have been higher. This emphasizes the role of the reverse shock and mantle binding energy in the fallback process whose effects are contingent upon the presence and structure of the hydrogen envelope. In any case, the explosion energy would be little changed and would be similarly large. Hence, we conclude that depending upon the degree of binary stripping (and likely the metallicity) this exemplar of black hole formation would result in a high explosion energy, but a spectrum of ejecta masses from a few to many $M_{\odot}$, an ejected 56Ni mass from high to lower, and a black hole mass from the lower gap to as much as perhaps $\sim$10.0 $M_{\odot}$. In all cases, a supernova explosion results.
We generally get a large spin parameter (as high as $\sim$0.6) that is very much a function of the character of the impact parameters of the (early)infall and (later)fallback plumes, themselves dependent on the chaotic flow field (Burrows et al., 2024b). The spin parameters of observed black-hole X-ray binaries (Miller et al., 2009) are $\sim$0.01--0.05. Hence, we are not sure how robust this number is for such a model. This is just want we find. If true, such a high value would suggest that this model, if properly simulated with B-fields, could be a collapsar model for long-soft gamma-ray bursts (Woosley, 1993). This is surprising since the initial model was non-rotating and the spin is induced (Rantsiou et al., 2011; Burrows et al., 2024b) during the early evolution.
@@
For the 19.56-$M_{\odot}$ model, the induced spin parameter ($a$) of the black hole at the time of its formation is only 0.006, the baryon mass (interior to 1011 g cm-3) is 2.278 $M_{\odot}$, and the corresponding gravitational mass is $\sim$2.071 $M_{\odot}$. At black hole formation, we mapped the Fornax model into FLASH and continued the calculation forward until $\sim$1.3$\times$105 seconds after bounce with a point mass and a diode inner boundary condition at 500 km. Stills of the subsequent evolution using FLASH are given in Figure 6. Note that by the end of this series the shock had traversed half of the progenitor’s radius and fallback has ceased.
@@
After the 19.56-$M_{\odot}$ model’s explosion has asymptoted, the gravitational mass of the black hole residue is $\sim$3.12 $M_{\odot}$, the ejected 56Ni mass is $\sim$0.20 $M_{\odot}$, the explosion energy is still $\sim$2.5 Bethes, and the total ejecta mass is $\sim$10.5 $M_{\odot}$. The net jet asymmetry results in a large (asymptotic) recoil kick to the PNS of $\sim$1300 km s-1, only $\sim$21 km s-1 of which derives from the net neutrino impulse (calculated at the PNS surface). Approximately $\sim$0.03 $M_{\odot}$ of 56Ni had fallen back. The lower residual gravitational mass for the 19.56-$M_{\odot}$ model at black hole formation and terminally than witnessed for the 40-$M_{\odot}$ model at the corresponding phases reflects 1) the longer time to black hole formation (and the resulting higher explosion energy), itself a reflection of the slightly lower post-bounce $\dot{M}$s and (crudely) compactness for the 19.56-$M_{\odot}$ model; and 2) the larger mantle binding energy of the 40-$M_{\odot}$ model. Importantly, as noted earlier, due to the larger mantle binding energy for the 40-$M_{\odot}$ model, it experiences significantly more fallback than the 19.56-$M_{\odot}$ model. The final black hole mass of the 19.56-$M_{\odot}$ model is clearly in the lower gap (similar to GW230529). Given its high kick speed, and depending upon the birth statistics, this subset of Channel 1 black holes would help make the observed lower mass gap in fact a true observational gap.
@@
The lower early induced spin for the 19.56-$M_{\odot}$ model seems related to the lower average impact parameters, which themselves may be slightly random. Again, an understanding of this difference awaits a more complete suite of 3D simulations to tease out the associated distributions that emerge from such chaotic behavior.
@@
Therefore, we suggest that Channel 1 objects leave behind black holes with masses both in and above the lower mass gap, with respectable supernova explosion energies, significant 56Ni yields, a wide range of ejecta masses, and a spectrum of kicks from modest to large. Recall that these specific quantitative outcomes depend upon the nuclear equation of state employed. We surmise that a nuclear equation of state with a higher maximum cold neutron star mass would result in a more energetic explosion and a lower final black hole mass. However, this too remains to be demonstrated.
@@
Figure 7: Sample snapshots of x-y slices in entropy space (per baryon per Boltzmann’s constant) of the later explosion of the 23-$M_{\odot}$ solar-metallicity model for six different times after the model was mapped from Fornax (top two) to FLASH (bottom four). Within $\sim$1000 seconds after bounce, the mass cut between the ejecta and the fallback matter destined to be incorporated into the black hole was determined. A total of $\sim$4.9 $M_{\odot}$ is incorporated into the black hole, while $\sim$9.9 $M_{\odot}$ is ejected in a modest supernova explosion with an energy of $\sim$0.45 Bethes. Note the large changes in scale from still to still among the last four panels, with the corresponding change in color maps and time after bounce (1, 7, 14, 68, 5336, and 310084 seconds). The last still is after the shock breakout from the progenitor star and the onset of Rayleigh-Taylor-like instabilities at the hydrogen/helium interface. See text for a discussion.
@@ 4.2 Channel 2: Black Hole Formation in the Context of Fallback and a Supernova Explosion
Interestingly, of the 0.088 $M_{\odot}$ of 56Ni that was explosively generated in this 23-$M_{\odot}$ model (Wang & Burrows, 2024b; Burrows et al., 2024a), $\sim$0.039 $M_{\odot}$ was eventually ejected. This was due to the outer fraction of the 56Ni pockets that acted like bullets and decelerated less than the inner fraction, thereby penetrating into the stellar mantle. Moreover, though we had originally determined that this model had a kick of $\sim$280 km s-1 (Burrows et al., 2024b), with fallback the net kick is partially due to the neutrino component, which then translates for this model into a black hole kick of $\sim$90 km s-1. However, there is no reason in principle that the net kick couldn’t be larger or smaller.
Therefore, what we find is a model that creates a $\sim$4.9 $M_{\odot}$ black hole, but explodes like a regular supernova with a modest explosion energy and 56Ni yield and a low recoil kick speed. This is not a “silent” supernova! For slightly greater explosion energies and/or less bound envelopes, the residual black hole could have been well into the lower mass gap and the kicks could have been higher. This variation might arise from a different CCSN explosion code and different massive-star progenitor structures. The converse is also true, wherein a larger mass black hole, accompanied by a still weaker explosion, might result. Given this, we can speculate that this channel of black hole formation might yield black holes with masses from the lower mass gap to $\sim$10 $M_{\odot}$. Importantly, a set of similar exploding fallback scenarios was explored by Chan et al. (2020).
As we noted in §4.1, if this model had experienced binary stripping, the residual black hole mass, ejected mass, and 56Ni yield could have been different. Indeed, with significant stripping this same model might have given birth to a neutron star instead of a black hole, emphasizing the potential importance of binary interaction in determining the outcome of collapse. We emphasize that our numbers arise from the marriage of our supernova code with KEPLER progenitors at solar-metallicity. Changing the 3D CCSN code, stellar evolution module, mass loss prescription (for a singlet or a binary), and metallicity can change the numbers. However, we suggest that the existence of this channel of black hole formation with modest energy and low kicks may well be robust. It is merely the mapping of progenitor ZAMS mass and metallicity to black hole mass, explosion energy, kick speed, and 56Ni yield that needs honing.
@@
Figure 9: Sample snapshots of x-y slices in entropy space (per baryon per Boltzmann’s constant) of the early explosion of the tenth-solar-metallicity 100-$M_{\odot}$ for four different times after bounce. Red is higher entropy and blue is the lower entropy of the progenitor matter into which the blast initially runs. The horizontal scale is 22,000 km. The upper left panel depicts the doomed explosion just $\sim$50 milliseconds after the black hole forms, which happens at $\sim$0.44 seconds after bounce. The dark-blue/light-blue transition in the top left and top right panels is at the oxygen/carbon interface. At the time of the top right panel the shock wave has penetrated this interface and for the bottom two panels it has left the graphical domain. After $\sim$10 seconds (not shown) the shock has weakened into a sound wave. As the bottom two panels suggest, within mere seconds the ejecta start to stall and fallback. Within $\sim$50 seconds, a $\sim$37 $M_{\odot}$ black hole near the putative lower boundary of the upper black hole mass gap has been created in the interior. See text for a discussion.
@@ 4.3 Channel 3: A Pulsational-Pair-Instability Progenitor: 100 Solar Masses
At very high stellar ZAMS masses, if high helium and CO core masses remain at the terminal stage of the star (generally thought to be associated with low wind mass loss rates and low metallicities), the envelope of a massive star is sufficiently pair- and radiation-dominated that it experiences an envelope instability and episodic thermonuclear events and mass loss in what is called a pulsational-pair instability supernova (PPISN) (Barkat et al., 1967; Woosley et al., 2002, 2007; Renzo & Smith, 2024). The core, however, of such a star still experiences core collapse via the Chandrasekhar instability and it is suggested that a stellar mass black hole would result. At even higher initial stellar and helium-core masses, stars are thought to detonate completely in a pair-instability supernova (PISN) (Fowler & Hoyle, 1964; Barkat et al., 1967) and leave no residue. Therefore, the maximum mass black hole formed via the PPISN is said to cap the mass of a black hole that can be birthed at the terminal stages of single massive star evolution, above which there is posited to be a black hole mass gap. As indicated earlier, this is but one of the gaps discussed in the context of the overall black hole mass spectrum  [^9] As noted earlier, the other (lower) mass gap discussed is between the maximum mass of a neutron star and the mass of the lightest measured black hole (near $\sim$5 $M_{\odot}$; MacLeod & Grindlay (2023)). Whether this is observationally a completely empty gap is the subject of current research. and could be near $\sim$40 $M_{\odot}$. However, there is a strong metallicity dependence to this bound (Belczynski et al., 2002, 2010) and its theoretical value is still in play. Nevertheless, PPISNe are one modality for stellar-mass black hole formation and have a bearing on the interpretation of the spectrum of black holes masses inferred by the LVK consortium detecting merging black hole binaries via gravitational radiation (Abbott et al., 2016; Aasi et al., 2015; Acernese et al., 2015; Akutsu et al., 2021).
@@
The interval between explosion ($\sim$250 milliseconds) and black hole formation ($\sim$440 milliseconds) for the 100 $M_{\odot}$ model was less than $\sim$200 milliseconds. This contrasts with the $\sim$1.5-second duration for the 40-$M_{\odot}$ model (Burrows et al., 2023) and the $\sim$3.5 duration for the 19.56 $M_{\odot}$ model (§4.1). In both the latter cases, there was sufficient time for the high neutrino absorption rate associated with the high post-explosion infall accretion luminosity (due to the high compactness in their stellar mantles and the breaking of spherical symmetry) to unbind some or most of the residual mass of the progenitor star. The result in those cases, discussed in §4.1, was an energetic asymmeterical supernova explosion with a high recoil kick and residual black holes either in or above the lower mass gap.

### v2

@@ 4 Black Hole Formation Channels
The different primary stellar-mass black hole formation pathways we discuss in this paper are represented by four distinct channels. For each channel we provide example models: Channel 1: asymmetric, high-compactness vigorously exploding solar-metallicity 40-$M_{\odot}$ and 19.56-$M_{\odot}$ models (§4.1) that within seconds of bounce form black holes, but later due to subsequent fallback may leave a spectrum of black hole masses; Channel 2: a weakly exploding, mildly asymmetric, solar-metallicity 23-$M_{\odot}$ model (§4.2) that experiences late-time fallback accretion and black hole formation on $\sim$1000-second timescales.; Channel 3: a 100-$M_{\odot}$ tenth-solar model (§4.3) that experiences a PPISN and then an aborted highly-asymmetrical supernova explosion that initially sends out a shock wave to thousands of kilometers which weakens significantly into a sound wave. Within hundreds of milliseconds, its PNS collapses into a black hole and within tens of seconds that black hole fattens to $\sim$37 $M_{\odot}$; and Channel 4: two models (12.25- and 14-$M_{\odot}$) (§4.4) that never reignited their stalled shock wave and would accrete over minutes to hours timescales until “quiescent” black hole formation. This is the oft-assumed “silent” channel. That this progenitor mass interval might evolve through this channel is quite unexpected and may be wrong. However, using the Sukhbold et al. (2018) progenitor suite, this is what we and others (Couch et al., 2020; Pan et al., 2021; Sykes & Müller, 2024b) currently conclude. We do not discuss the secondary black hole formation channels through accretion onto a neutron star from a companion star, the merger of two neutron stars, or the late-time developments of a Thorne-Zytkov object (Thorne & Zytkow, 1977), all potentially important pathways, but outside the CCSN context.
Figure 5: Snapshots of an x-y slice in entropy space (per baryon per Boltzmann’s constant) of the inner region ($\pm$1000 km) of the explosion of the solar-metallicity 19.56-$M_{\odot}$ progenitor. These stills are taken at 1.0, 2.0, 3.0, and 3.767 seconds after bounce, the latter just before black hole formation. They reveal an asymmetric quasi-jetlike structure flapping about in response to the variations in the infalling plumes that are dancing over the PNS. The red regions are high-entropy neutrino-driven ejected, while the blue regions are mostly these lower-entropy (and more dense) infalling streams. Note the blue and red regions are “complementary” -- the driven ejecta avoid the infalling fingers. The degree of net asymmetry is correlated with the magnitude of the resultant recoil kick, with the corresponding neutrino emission anisotropy a small contributing factor. The 19.56-$M_{\odot}$ model explodes vigorously and asymmetrically and leaves behind a black hole. See text for discussion.
@@
Interestingly, if this 40-$M_{\odot}$ model had experienced binary mass transfer and been stripped to a “stripped-envelope” structure, the fallback mass would have been minimal, the ejected 56Ni mass would have been more significant, the total ejected mass would have been much lower, as would the residual black hole mass, and the kick to the black hole would have been higher. This emphasizes the role of the reverse shock and mantle binding energy in the fallback process whose effects are contingent upon the presence and structure of the hydrogen envelope. In any case, the explosion energy would be little changed and would be similarly large. Hence, we speculate that depending upon the degree of binary stripping (and likely the metallicity) this exemplar of black hole formation would result in a high explosion energy, but a spectrum of ejecta masses from a few to many $M_{\odot}$, an ejected 56Ni mass from high to lower, and a black hole masses up to perhaps $\sim$10.0 $M_{\odot}$. In all cases, a supernova explosion results.
We generally get a large spin parameter (as high as $\sim$0.6) [^10] The spin parameter ($a$) is defined as $\frac{L}{GM^{2}/c}$, where $L$ is the total angular momentum and $M$ is the core gravitational mass, and the other quantities have their usual meanings. The maximum possible $a$ is 1.0. that is very much a function of the character of the impact parameters of the (early)infall and (later)fallback plumes, themselves dependent on the chaotic flow field (Burrows et al., 2024b). The spin parameters of observed black-hole X-ray binaries (Miller et al., 2009) are $\sim$0.01--0.05. Hence, we are not sure how robust this number is for such a model. This is just want we find. If true, such a high value would suggest that this model, if properly simulated with B-fields, could be a collapsar model for long-soft gamma-ray bursts (Woosley, 1993). This is surprising since the initial model was non-rotating and the spin is induced (Rantsiou et al., 2011; Burrows et al., 2024b) during the early evolution.
@@
For the 19.56-$M_{\odot}$ model, the induced spin parameter ($a$) of the black hole at the time of its formation is only 0.006 (and at the end of the FLASH phase of the simulation is $\sim$0.09), the baryon mass (interior to 1011 g cm-3) is 2.278 $M_{\odot}$, and the corresponding gravitational mass is $\sim$2.071 $M_{\odot}$. At black hole formation, we mapped the Fornax model into FLASH and continued the calculation forward until $\sim$1.3$\times$105 seconds after bounce with a point mass and a diode inner boundary condition at 500 km. Stills of the subsequent evolution using FLASH are given in Figure 7. Note that by the end of this series the shock had traversed half of the progenitor’s radius and fallback has ceased.
@@
After the 19.56-$M_{\odot}$ model’s explosion has asymptoted, the gravitational mass of the black hole residue is $\sim$3.12 $M_{\odot}$, the ejected 56Ni mass is $\sim$0.20 $M_{\odot}$, the explosion energy is still $\sim$2.5 Bethes, and the total ejecta mass is $\sim$10.5 $M_{\odot}$. The net jet asymmetry results in a large (asymptotic) recoil kick to the PNS of $\sim$1300 km s-1, only $\sim$21 km s-1 of which derives from the net neutrino impulse (calculated at the PNS surface). Approximately $\sim$0.03 $M_{\odot}$ of 56Ni had fallen back. The lower residual gravitational mass for the 19.56-$M_{\odot}$ model at black hole formation and terminally than witnessed for the 40-$M_{\odot}$ model at the corresponding phases reflects 1) the longer time to black hole formation (and the resulting higher explosion energy), itself a reflection of the slightly lower post-bounce $\dot{M}$s and (crudely) compactness for the 19.56-$M_{\odot}$ model; and 2) the larger mantle binding energy of the 40-$M_{\odot}$ model. Importantly, as noted earlier, due to the larger mantle binding energy for the 40-$M_{\odot}$ model, it experiences significantly more fallback than the 19.56-$M_{\odot}$ model. The final black hole mass of the 19.56-$M_{\odot}$ model is clearly in the putative lower black hole mass gap (Shao, 2022) (similar to GW230529). Given its high kick speed, and depending upon the birth statistics, this subset of Channel 1 black holes would help make the observed lower mass gap in fact a true observational gap.
@@
The lower final induced spin parameter for the 19.56-$M_{\odot}$ model ($\sim$0.09, as stated above) seems related to the lower average impact parameters, which themselves may be slightly random. Again, an understanding of this difference awaits a more complete suite of 3D simulations to tease out the associated distributions that emerge from such chaotic behavior.
@@
Therefore, we suggest that Channel 1 objects leave behind black holes, and result in respectable supernova explosion energies, significant 56Ni yields, a wide range of possible ejecta masses, and a spectrum of kicks from modest to large. Recall that these specific quantitative outcomes depend upon the nuclear equation of state employed. We surmise that a nuclear equation of state with a higher maximum cold neutron star mass would result in a more energetic explosion and a lower final black hole mass. However, this too remains to be demonstrated.
@@
Figure 8: Sample snapshots of x-y slices in entropy space (per baryon per Boltzmann’s constant) of the later explosion of the 23-$M_{\odot}$ solar-metallicity model for six different times after the model was mapped from Fornax (top two) to FLASH (bottom four). Within $\sim$1000 seconds after bounce, the mass cut between the ejecta and the fallback matter destined to be incorporated into the black hole was determined. A total of $\sim$4.9 $M_{\odot}$ is incorporated into the black hole, while $\sim$9.9 $M_{\odot}$ is ejected in a supernova explosion with an energy of $\sim$0.45 Bethes. Note the large changes in scale from still to still among the last four panels, with the corresponding change in color maps and time after bounce (1, 7, 14, 68, 5336, and 310084 seconds). The last still is after the shock breakout from the progenitor star and the onset of Rayleigh-Taylor-like instabilities at the hydrogen/helium interface. See text for a discussion.
@@ 4.2 Channel 2: Black Hole Formation in the Context of Fallback and a Supernova Explosion
Interestingly, of the 0.088 $M_{\odot}$ of 56Ni that was explosively generated in this 23-$M_{\odot}$ model (Wang & Burrows, 2024c; Burrows et al., 2024a), $\sim$0.039 $M_{\odot}$ was eventually ejected. This was due to the outer fraction of the 56Ni pockets that acted like bullets and decelerated less than the inner fraction, thereby penetrating into the stellar mantle. Moreover, though we had originally determined that this model had a kick of $\sim$280 km s-1 (Burrows et al., 2024b), with fallback the net kick is partially due to the neutrino component, which then translates for this model into a black hole kick of $\sim$90 km s-1. However, there is no reason in principle that the net kick couldn’t be larger or smaller. The corresponding final spin parameter ($a$) for this model is $\sim$0.135.
Therefore, what we find is a model that creates a $\sim$4.9 $M_{\odot}$ black hole, but explodes like a regular supernova with a modest (by which we mean in the mid-range of inferred measured values and our theoretical results) explosion energy and 56Ni yield and a low recoil kick speed. This is not a “silent” supernova! For slightly greater explosion energies and/or less bound envelopes, the residual black hole could have been lighter and the kicks could have been higher. This variation might arise from a different CCSN explosion code and different massive-star progenitor structures. The converse is also true, wherein a larger mass black hole, accompanied by a still weaker explosion, might result. Given this, we can speculate that this channel of black hole formation might yield black holes with a range of masses, yet to be determined. Importantly, a set of similar exploding fallback scenarios was explored by Chan et al. (2020).
As we noted in §4.1, if this model had experienced binary stripping, the residual black hole mass, ejected mass, and 56Ni yield could have been different. Indeed, with significant stripping this same model might have given birth to a neutron star instead of a black hole, emphasizing the potential importance of binary interaction in determining the outcome of collapse. We emphasize that our numbers arise from the marriage of our supernova code with KEPLER progenitors at solar-metallicity. Changing the 3D CCSN code, stellar evolution module, mass loss prescription (for a singlet or a binary), and metallicity can change the numbers. However, we suggest that the existence of this channel of black hole formation with intermediate energy and low kicks may well be robust. It is merely the mapping of progenitor ZAMS mass and metallicity to black hole mass, explosion energy, kick speed, and 56Ni yield that needs honing.
@@
Figure 10: Sample snapshots of x-y slices in entropy space (per baryon per Boltzmann’s constant) of the early explosion of the tenth-solar-metallicity 100-$M_{\odot}$ for four different times after bounce. Red is higher entropy and blue is the lower entropy of the progenitor matter into which the blast initially runs. The horizontal scale is 22,000 km. The upper left panel depicts the doomed explosion just $\sim$50 milliseconds after the black hole forms, which happens at $\sim$0.44 seconds after bounce. The dark-blue/light-blue transition in the top left and top right panels is at the oxygen/carbon interface. At the time of the top right panel the shock wave has penetrated this interface and for the bottom two panels it has left the graphical domain. After $\sim$10 seconds (not shown) the shock has weakened into a sound wave. As the bottom two panels suggest, within mere seconds the ejecta start to stall and fallback. Within $\sim$50 seconds, a $\sim$37 $M_{\odot}$ black hole has been created in the interior. See text for a discussion.
@@ 4.3 Channel 3: A Pulsational-Pair-Instability Progenitor: 100 Solar Masses
The interval between explosion ($\sim$250 milliseconds) and black hole formation ($\sim$440 milliseconds) for the 100 $M_{\odot}$ model was less than $\sim$200 milliseconds. This contrasts with the $\sim$1.5-second duration for the 40-$M_{\odot}$ model (Burrows et al., 2023) and the $\sim$3.5 duration for the 19.56 $M_{\odot}$ model (§4.1). In both the latter cases, there was sufficient time for the high neutrino absorption rate associated with the high post-explosion infall accretion luminosity (due to the high compactness in their stellar mantles and the breaking of spherical symmetry) to unbind some or most of the residual mass of the progenitor star. The result in those cases, discussed in §4.1, was an energetic asymmeterical supernova explosion with a black hole residue and a high recoil kick.